# Multi-Keypoint Affordance Representation for Functional Dexterous Grasping

### v1

@@ III-A2 Keypoint Feature Extraction from Egocentric View
To extract keypoint features from the Ego view image, we define a set of learnable weights $W\in\mathbb{R}^{t\times n}$, where $t$ represents the number of task types and $n$ is the number of candidate keypoints. These weights are multiplied with the candidate keypoints $K_{cn}$ to select the three keypoints $K_{i}$ (where $i=1,2,3$) for feature extraction from the corresponding regions in the Ego view image.
@@
For the selected keypoints $K_{i}$, we define a circular region centered at each keypoint with a radius $r$ and extract features from these regions. The extracted region features are denoted as $F_{ki}$, representing the features from the regions centered on the selected keypoints.
To align the features from the Ego and Exo views in a unified feature space, we apply a linear transformation to the extracted keypoint features $F_{ki}$, resulting in the final keypoint features $f_{op}$:
$f_{ki}=\text{proj}(F_{ki}),$
where the projection layer proj is a linear transformation that maps the Ego view features to the feature space aligned with the Exo view.
@@ III-A3 Contact Geometry Knowledge Transfer
In the final step, we utilize the CAM technique [16] to classify Exo’s features for the specific task, with the classification loss denoted as $L_{\text{cls}}$. Additionally, we extract object features from the interactive regions using the Extract and PartSelect [38] modules, obtaining the prototype features $f_{\text{op}}$ for the keypoint regions.
For the three keypoint features $f_{i}$ extracted from the Ego view, we compute their sum to obtain the global keypoint feature $f_{\text{gk}}$, which encapsulates the contact geometry information from the Ego perspective:
$f_{\text{gk}}=\sum_{i=1}^{3}f_{ki}.$
@@
$L_{\text{cos}}=1-\frac{f_{\text{op}}\cdot f_{\text{gk}}}{\|f_{\text{op}}\|\|f_{\text{gk}}\|}.$
@@ III-B Keypoint-based Grasp Matrix Transformation
$R_{O}^{I}=[\mathbf{x}_{o},\mathbf{y}_{o},\mathbf{z}_{o}].$
At the same time, obtain the transformation matrix between the world coordinate system $I$ and the object coordinate system $O$:
$T_{O}^{I}=\begin{bmatrix}R_{O}^{I}&W_{o}\\ 0&1\end{bmatrix}$
@@
$R_{H}^{I}=[\mathbf{x}_{h},\mathbf{y}_{h},\mathbf{z}_{h}].$
@@ IV Experiments
IV-A Setup
Datasets: The public challenging FAH benchmark [17] is a large-scale affordance-annotated dataset specifically designed for hand-object interactions. It contains $6$ functional grasp affordances and $18$ tools, with $,85$ 858 images spanning both exocentric (Exo) and egocentric (Ego) views. The dataset provides image-level affordance labels for weakly supervised learning and annotations for coarse dexterous grasp gestures targeting specific “Task Tool” pairs. However, its test set only includes heatmap annotations for functional finger contact regions. To address this limitation, we sparsely annotate two additional contact points (little finger and wrist projection points). Specifically, polygons with up to five points are constructed around finger keypoints within a $5mm$ radius, and Gaussian kernels are applied at each point to generate dense annotations. During training, point annotations are added to the object regions in each Ego-view image to distinguish foreground and background during segmentation with SAM [14].
Implementation Details: Experiments are conducted on an NVIDIA RTX A6000 GPU. The model is trained using the SGD optimizer with a learning rate of $0.01$ over $15$ iterations. Images are resized to a resolution of $448{\times}448$. Following prior works [38, 4], we evaluate the results using Kullback-Leibler Divergence (KLD), Similarity (SIM), and Normalized Scanpath Saliency (NSS) metrics.
@@ IV-B Results of Functional Affordance Grounding
In this section, we present qualitative and quantitative results to demonstrate the effectiveness and efficiency of our method on the FAH test set [17]. As weakly- or unsupervised methods for multi-region affordance localization are scarce in the state of the art, we use ReKep* [13], a keypoint prediction method, as our baseline.
@@
Hyperparameter Analysis. We further investigate the impact of the candidate keypoint number $N{=}\{6,9,12,15\}$ on model performance, In Tab. II, we show the effects of different values on KLD, SIM, and NSS. The results indicate that $N{=}12$ achieves the best performance across all metrics. This aligns with our design principle: too few keypoints lead to insufficient representation of affordance regions, hindering fine-grained grasp modeling, while too many introduce redundancy, diluting feature importance and reducing the model’s focus on functional regions.
Ablation Study. The object priors provided by SAM [14] are crucial for constraining keypoint proposals to objects in the scene rather than the background. Thus, we focus on analyzing the critical visual feature extraction network in CMKA. As shown in Tab. III, DINOv2 [15], as the backbone network combined with our designed multi-scale feature fusion (MSFF) module, achieves the best performance. In the backbone network, DINOv2 generates clearer features compared to DINO-ViT [41], better distinguishing fine-grained object regions and leading to improved performance. Furthermore, compared to replacing MSFF with a simple fully connected network, MSFF, with its multi-layer and multi-scale feature mapping, demonstrates superior potential.
TABLE II: Impact of the candidate keypoint number $N$.
$N$ KLD ($\downarrow$) SIM ($\uparrow$) NSS ($\uparrow$)
6 5.409 0.308 0.849
9 5.748 0.282 0.737
12 5.035 0.313 0.865
15 5.766 0.279 0.723
TABLE III: Ablation study on different feature extractors. FFL: feed-forward layer. MSFF: multi-scale feature fusion.
DINOv2 DINO-ViT MSFF FFL KLD ($\downarrow$) SIM ($\uparrow$) NSS ($\uparrow$)
@@
Qualitative Analysis. As shown in Fig. 6, the visibility grounding visualizations include Ground Truth (GT), our method, and the baseline method ReKep* [13]. Compared to the baseline, our method localizes keypoints within the hand-object contact region while preserving the relative spatial relationships among the functional finger, little finger, and wrist projection, ensuring a meaningful distribution for dexterous grasping. For example, in the “Click Flashlight” task, keypoint 0 is localized on the thumb and keypoint 1 on the little finger, accurately reflecting the contact regions. In the “Press Drill” task, keypoint 0 is placed on the index finger, keypoint 1 on the little finger, and keypoint 2 on the wrist projection, maintaining a reasonable spatial relationship. In contrast, ReKep* [13] relies on manual post-processing, failing to constrain keypoints within the contact region and lacking spatial consistency, resulting in scattered keypoints and reduced accuracy of the affordance region.
@@ IV-D Performance in Real-World Scenarios
Figure 8: Experimental setup in a real-world scenario: (a) Hardware platform; (b) Calibration standards for the geometric relationship between keypoints when the functional finger is the index finger (upper) and the thumb (lower).
Figure 9: Experiments with three typical “Task Tools” in real-world scenarios: “Click Flashlight”, “Press Drill” and “Press Spraybottle” (from top to bottom). The upper right corner of the fourth column compares the functional dexterous grasping success rate with our method and GAAF-Dex [17] (in bracket), where the total grasp number of each instance is $5$.
Real-world Experiments Setup: As shown in Fig. 8(a), the real-world platform consists of an Inspire hand, a UR5 industrial robotic arm, an Intel RealSense camera, a tool rack, and a control computer. To address real-world uncertainties, we introduce keypoint relative position calibration annotations based on the Inspire model during the grasping process, as shown in Fig. 8(b).
In the experiments, we use tool instances commonly found in daily life, which were unseen in the training set. As shown in Fig. 9, we selected three “Task Tool” with strict functional grasping requirements for the experiment: “Click Flashlight”, “Press Drill”, and “Press Spraybottle”. We recorded four states: the initial state, followed by the localization of three affordance keypoints on the tool surface using the CMKA method to estimate their initial planar relationship. Simultaneously, we utilized an RGB-D camera to obtain the depth $z$ from the $(x,y)$ coordinates of the keypoints in the depth map, thus acquiring their $(x,y,z)$ coordinates in 3D space. We then adjusted the relative geometric positions of the keypoints on the dexterous hand palm using calibration standards (yellow triangles in the third column). Finally, we apply the KGT method to compute the final wrist grasp pose $W(R,T)$, and use the coarse gesture labels from the FAH [17] to obtain the coarse grasp angle $J$ for each “Task Tool," which controls the final grasping of the dexterous hand.
The results demonstrate that for complex tasks requiring precise finger-object alignment, such as Press and Click, our method effectively bridges the gap between multi-point perception and dexterous grasping, showing broad practical value. Furthermore, due to the lack of direct methods combining perception and dexterous grasping, we compared our method with the state-of-the-art GAAF-Dex [17] by the number of successful functional grasps. We define a successful grasp as the functional finger combining with the tool’s functional component while the remaining fingers securely grasp other parts of the tool. As shown in the top left corner of Fig. 9, our method improves the grasp success rate by an average of $40\%$ across three complex tasks. GAAF-Dex [17] is only effective when the initial and final grasp rotations are similar, such as in the Click Flashlight scenario, due to its lack of adaptive rotation handling. In contrast, our method can handle arbitrary initial grasp poses.
@@ V Conclusion and Future Work
This work proposes a keypoint-based affordance representation for functional dexterous grasping. By leveraging human experience data for weak supervision and integrating the CMKA module with large visual models, our approach achieves precise multi-point contact localization, reducing data annotation costs and improving generalization. The KGT method enables the mapping of dexterous hand postures to object keypoints, ensuring a direct connection between perception and action. Experimental results demonstrate that the proposed method outperforms existing approaches in localization accuracy and functional grasp success rate. Practical experiments show that relying solely on 2D vision for localization fails to provide stable grasp constraints. In the future, we aim to utilize multimodal information to enhance the accuracy and stability of multi-point 3D localization in real-world scenarios.

### v2

@@ III-A2 Affordance-aware Keypoint Feature Extraction from Ego View
To extract keypoint features from the Ego view image, we define a set of learnable weights $W{\in}\mathbb{R}^{T\times N}$, where $T$ represents the number of affordance classes and $N$ is the number of candidate keypoints. These weights are multiplied with the candidate keypoints $K_{n}$ to select the final three keypoints $K_{i}$ (where i=1,2,3i{=}1,2,3) for feature extraction from the corresponding regions in the Ego view image.
@@
For the selected keypoints $K_{i}$, we define a circular region centered at each keypoint with a radius $r$ and extract features from these regions, denoted as $F_{ki}$.
To align the features from the Ego and Exo views in a unified feature space, we project the features of the selected three keypoints and sum them to obtain the final keypoint feature $f_{gk}$, which encapsulates the contact geometry information from the Ego view:
$f_{gk}=\sum_{i=1}^{3}\text{proj}(F_{ki}),$ (5)
where $\text{proj}(\cdot)$ denotes a linear projection layer.
@@ III-A3 Contact Geometry Knowledge Transfer
To supervise egocentric keypoints selection, we define an affordance-specific activation function $\varphi_{\text{cam}}$ [14], which consists of a feed-forward layer, two convolutional layers, and a $1{\times}1$ class-aware convolution. Given the exocentric feature map $F_{exo}$, $\varphi_{\text{cam}}$ produces the affordance localization map $P{\in}\mathbb{R}^{C\times H\times W}$, where $C$ is the number of affordance classes. We apply global average pooling on $P$ to obtain class-wise affordance scores, which are optimized using a cross-entropy loss $L_{\text{cls}}$.
In addition, we adopt the Extract and PartSelect modules from LOCATE [38] to obtain the object-part prototype feature $f_{\text{op}}$, by filtering part embeddings through clustering and metric, resulting in representative contact features.
@@
$L_{\text{cos}}=1-\frac{f_{\text{op}}\cdot f_{\text{gk}}}{\|f_{\text{op}}\|\|f_{\text{gk}}\|}.$ (6)
@@ III-B Keypoint-based Grasp Matrix Transformation
$R_{O}^{I}=[\mathbf{x}_{o},\mathbf{y}_{o},\mathbf{z}_{o}].$ (8)
At the same time, we obtain the transformation matrix between the world coordinate system $I$ and the object coordinate system $O$:
$T_{O}^{I}=\begin{bmatrix}R_{O}^{I}&W_{o}\\ 0&1\end{bmatrix}$. (9)
@@
$R_{H}^{I}=[\mathbf{x}_{h},\mathbf{y}_{h},\mathbf{z}_{h}].$ (10)
@@ IV Experiments
IV-A Experiment Setup
Datasets: The public challenging FAH benchmark [15] is a large-scale affordance-annotated dataset specifically designed for hand-object interactions. It contains $6$ functional grasp affordances (such as “Press”, “Click”, “Hold”, etc.). and $18$ tools, with $5,858$ images spanning both Exo and Ego views. The dataset provides image-level affordance labels for weakly supervised learning and annotations for coarse dexterous grasp gestures targeting specific “Task Tool” pairs. However, its test set only includes heatmap annotations for functional finger contact regions. To address this limitation, we annotate two additional contact points (little finger and wrist projection areas). Specifically, polygons with up to five points are constructed around finger keypoints within a radius of $5mm$, and Gaussian kernels are applied at each point to generate dense annotations. During training, point annotations are added to the object regions in each Ego-view image to distinguish foreground and background during segmentation with SAM [32].
Implementation Details: Experiments are conducted on an NVIDIA RTX A6000 GPU. The model is trained using the SGD optimizer with a learning rate of $0.01$ over $15$ iterations. Images are resized to a resolution of $448{\times}448$.
Metrics: For affordance grounding, we adopt Kullback-Leibler Divergence (KLD), Similarity (SIM), and Normalized Scanpath Saliency (NSS) following prior work [4, 38].
For functional grasping, we measure the functional grasp success rate (FGS) as defined in [25]: a grasp is successful if the hand holds the object stably for at least ten seconds and correctly performs the intended action on the tool’s functional area.
In addition, we introduce the 2D-to-Physical Contact Consistency (TPC) metric to evaluate whether the predicted 2D keypoints, when projected into 3D, fall within the functional contact region. TPC is defined as the ratio of the number of hits $n$ to the three selected keypoints.
@@ IV-B Results of Functional Affordance Grounding
In this section, we present qualitative and quantitative results to demonstrate the effectiveness of our method on the FAH test set [15]. As weakly supervised methods for multi-region affordance localization are scarce in the state of the art, we use ReKep* [13], a keypoint prediction method, as our baseline.
@@
Hyperparameter Sensitivity Analysis. To evaluate the impact of the total number of candidate keypoints $N{=}S{\times}J$ on model performance, we systematically test different combinations of region numbers $S{\in}\{2,3,4\}$ and cluster centers per region $J{\in}\{2,3,4,5\}$, using KLD, SIM, and NSS as evaluation metrics (see Fig. 7). Experimental results show that the configuration S=3S{=}3, J=4J{=}4 (i.e., N=12N{=}12) achieves the best performance across all three metrics.
In general, moderately increasing the number of candidate keypoints helps improve spatial coverage and semantic diversity, enabling the model to better perceive task-relevant regions. However, an excessive number of keypoints may introduce redundancy and noise, thereby hindering the learning of discriminative weights. Notably, under the condition S=3S{=}3, the performance of J=3J{=}3 is not only lower than that of J=4J{=}4, but also slightly worse than J=2J{=}2. This indicates that under weak supervision, the model benefits more from structurally coherent and well-distributed candidate arrangements rather than simply increasing the number of keypoints.
Ablation Study. The object priors provided by SAM [32] are crucial for constraining keypoint proposals to objects in the scene rather than the background. Thus, we focus on analyzing the critical visual feature extraction network in CMKA. As shown in Tab. II, DINOv2 [31], as the backbone network combined with our designed Hierarchical Feature Fusion (HFF) module, achieves the best performance. In the backbone network, DINOv2 generates clearer features compared to DINO-ViT [41], better distinguishing fine-grained object regions and leading to improved performance. Furthermore, compared to replacing HFF with a simple fully connected network, HFF, with its multi-layer and multi-scale feature mapping, demonstrates superior potential.
Figure 7: Performance heatmaps for combinations of region number $S$ and cluster number $J$. The best results are consistently achieved at S=3S{=}3, J=4J{=}4 (marked by ★).
TABLE II: Ablation study on different feature extractors. FFL: feed-forward layer. HFF: hierarchical feature fusion.
DINOv2 DINO-ViT HFF FFL KLD ($\downarrow$) SIM ($\uparrow$) NSS ($\uparrow$)
@@
Qualitative Analysis. As shown in Fig. 6, we compare visibility grounding results from Ground Truth (GT), our method, and the baseline ReKep* [13]. Our method accurately localizes keypoints within the hand-object contact region while preserving the spatial relationships among the functional finger, little finger, and wrist, ensuring a meaningful distribution for dexterous grasping. For example, in “Click Flashlight”, keypoints correctly align with the thumb and little finger, while in “Press Drill”, they align with the index finger, little finger, and wrist. In contrast, ReKep* [13], which relies on manual post-processing, fails to constrain keypoints within the contact region and lacks spatial consistency, resulting in scattered and less reliable affordance localization.
@@ IV-D Performance in Real-World Scenarios
Figure 9: Experiments with three typical “Task Tool” in real-world scenarios. (a) Hardware setup. (b) Calibration standards illustrating the geometric relationship between keypoints when the functional finger is the index finger (up) and the thumb (down). (c) Functional grasping steps across the 1st to 4th rows: Initial, Localization, Adjustment, and Control.
As shown in Fig. 9 (a), the real-world platform consists of an Inspire hand, a UR5 industrial robotic arm, an Intel RealSense camera, a tool rack, and a control computer. To address real-world uncertainties, we introduce keypoint relative position calibration annotations based on the Inspire model during the grasping process, as shown in Fig. 9(b).
In the experiments, we evaluated common daily tools unseen during training, including three “Task-Tool” pairs with strict functional grasping requirements: “Click Flashlight”, “Press Drill”, and “Press Spraybottle” (Fig. 9(c)). The process includes four steps: (1) CMKA localizes three affordance keypoints and estimates their planar relationship; (2) depth values at the 2D locations are retrieved from the RGB-D depth map to obtain 3D coordinates (see the bottom of the second row); (3) keypoints are adjusted on the hand palm based on calibration (see the yellow triangles in the third row); (4) KGT computes the wrist grasp pose $W(R,T)$, combined with a coarse grasp angle $J$ from FAH [15], to execute the grasp. Results show that our method effectively bridges perception and execution, enabling precise finger-object alignment in complex tasks like “Press” and “Click”.
TABLE III: Results of FGS / TPC (%) on three representative “Task-Tool” in real-world scenarios.
Method Click Flashlight Press Drill Press Spraybottle
GAAF-Dex [15] 60 / – 0 / – 0 / –
Ours 80 / 66.7 60 / 100 40 / 66.7
Furthermore, due to the lack of direct methods combining perception and dexterous grasping, we compared our method with the state-of-the-art GAAF-Dex [15] by the functional grasp success rate (FGS). As shown in Tab. III, cross the three complex tasks, our method achieves an average FGS of 60%, compared to 20% for GAAF-Dex, corresponding to an absolute improvement of 40 percentage points. GAAF-Dex relies on the similarity between initial and target grasp orientations, succeeding only in cases like “Click Flashlight”, but failing in others due to the lack of rotation adaptation. In contrast, our method handles arbitrary initial poses.
Tab. III reports the 2D-to-Physical contact consistency (TPC) results. Our method achieves $100\%$ on “Press Drill” and $66.7\%$ on both “Click Flashlight” and “Press Spraybottle”, showing reliable contact alignment across tasks, though performance varies due to challenges like depth ambiguity and transparency. While high TPC does not ensure high FGS, the results demonstrate good generalization of our method and provide a solid basis for further improvement with feedback and online adjustment.
@@ V Conclusion and Future Work
This work proposes a keypoint-based affordance representation for functional dexterous grasping. By leveraging human experience data for weak supervision and integrating the CMKA module with large visual models, the proposed method achieves precise multi-point contact localization with pose constraints for functional grasping, effectively reducing annotation costs and improving generalization. The KGT method enables the mapping of dexterous hand postures to object keypoints, ensuring a direct connection between perception and action. Experimental results demonstrate that our method outperforms existing approaches in both localization accuracy and functional grasp success rate. Real-world experiments further show that, while 2D vision alone provides reliable perception, it may be insufficient to ensure robust grasp execution under real-world uncertainties.
In the future, we aim to utilize multimodal information to enhance the accuracy and stability of multi-point 3D localization in real-world scenarios.